Please help not sure how to go about answering this problem, here goes: A business man invest a total of $13900 into two businesses A and B at the simple interest rate of 14% per annum and 11% per annum respectively. If the total amount of simple interest for 2 years was $3508, what was the amount invested in business B

Found 2 solutions by rothauserc, ankor@dixie-net.com:
Answer by rothauserc(4718)

let x be amount invested in business A and y amount invested in business B, then
x + y = 13900
now I = p*r*t
3508 = (.14x + .11y) * 2
solve first equation for x
x = 13900 - y
substitute for y in second equation
3508 = (.14(13900-y) + .11y) * 2
divide both sides of = by 2
1754 = (1946 -.14y + .11y)
1754 = (1946 -.03y)
.03y = 192
y = 6400
$6400 invested in business B

Answer by ankor@dixie-net.com(22740)

A business man invest a total of $13900 into two businesses A and B at the simple interest rate of 14% per annum and 11% per annum respectively.
If the total amount of simple interest for 2 years was $3508, what was the amount invested in business B.
:
Let a = amt invested at 14% (.14)
Let b = amt invested at 11% (.11)
therefore
a + b = 13900
b = (13900-a); use this form for substitution
:
" If the total amount of simple interest for 2 years was $3508, "
therefore (remember it's 2yrs simple interest)
2(.14a + .11b) = 3508
Divide both sides by 2
.14a + .11b = 1754
Substitute (13900-a) for b
.14a + .11(13900-a) = 1754
.14a + 1529 - .11a = 1754
.14a - .11a = 1754 - 1529
.03a = 225
a = 225/.03
a = 7500 invested at 14%
;
"what was the amount invested in business B"
b = 13900 - 7500
b = $6400 invested at 11%
:
:
Check this for yourself, find the interest of each using I - PRT
See that they add up to 3508